A chemist has two large containers of sulfuric acid solution, with different concentrations of acid in each container. Blending 320 mL of the first solution and 585 mL of the second gives a mixture that is 11.84% acid, whereas blending 80 mL of the first mixed with 490 mL of the second gives a 6.51% acid mixture. What are the concentrations of sulfuric acid in the original containers? (Round your answer to the nearest whole number.)
Answer by richwmiller(17219)

320x+585y=11.84*(585+320),
80x+490y=.0651*(80+490)

80*x+490*y=37.107
320*x+585*y=47.952
x=0.0162626, y=0.0730735
x=1.62626, y=7.30735.
x=2 y=7